# Optimizing Calibration by Gaining Aware of Prediction Correctness

### v2

@@ Appendix A Experimental Setup
A.1 ImageNet Setup
Models. We employ the ImageNet models from the PyTorch Image Models (timm) library (Wightman, 2019), which offers models trained or fine-tuned on the ImageNet-1k training set (Deng et al., 2009). The models utilized in our paper are listed below:
@@ A.1 ImageNet Setup
ImageNet-A(dversarial) (Hendrycks et al., 2021b):
https://github.com/hendrycks/natural-adv-examples.
ImageNet-S(ketch) (Wang et al., 2019):
https://github.com/HaohanWang/ImageNet-Sketch.
ImageNet-R(endition) (Hendrycks et al., 2021a):
https://github.com/hendrycks/imagenet-r.
ImageNet-Blur (Hendrycks & Dietterich, 2019):
https://github.com/hendrycks/robustness.
ObjectNet (Barbu et al., 2019):
https://objectnet.dev/download.html.
A.2 CIFAR-$10$ Setup
Models. We employ the CIFAR-$10$ models from the open source library (https://github.com/kuangliu/pytorch-cifar) which offers models trained or fine-tuned on the CIFAR-$10$ training set (Krizhevsky et al., 2009). The models utilized in our paper are listed below:
@@ A.2 CIFAR-$10$ Setup
https://github.com/BayesWatch/cinic-10.
CIFAR10-C (Hendrycks & Dietterich, 2019)(https://github.com/hendrycks/robustness);
A.3 iWildCam Setup
Models. We employ the iWildCam models from the open source library (https://worksheets.codalab.org/worksheets/ 0x52cea64d1d3f4fa89de326b4e31aa50a) which offers models trained or fine-tuned on the iWildCam training set (Beery et al., 2020). The models utilized in our paper are listed below:
@@
Appendix B Additional Visualisations
B.1 AUROC curves

### v3

@@ Appendix E Access Of Benchmarks And Models
ImageNet models. We employ the ImageNet models from the PyTorch Image Models (timm) library (Wightman, 2019), which offers models trained or fine-tuned on the ImageNet-1k training set (Deng et al., 2009). The models utilized in our paper are listed below:
@@
ImageNet-A(dversarial) (Hendrycks et al., 2021b): https://github.com/hendrycks/natural-adv-examples.
ImageNet-S(ketch) (Wang et al., 2019): https://github.com/HaohanWang/ImageNet-Sketch.
ImageNet-R(endition) (Hendrycks et al., 2021a): https://github.com/hendrycks/imagenet-r.
ImageNet-Blur (Hendrycks & Dietterich, 2019): https://github.com/hendrycks/robustness.
ObjectNet (Barbu et al., 2019): https://objectnet.dev/download.html.
CIFAR-$10$ models. We employ the CIFAR-$10$ models from the open source library (https://github.com/kuangliu/pytorch-cifar) which offers models trained or fine-tuned on the CIFAR-$10$ training set (Krizhevsky et al., 2009). The models utilized in our paper are listed below:
@@
https://github.com/BayesWatch/cinic-10. CIFAR10-C (Hendrycks & Dietterich, 2019)(https://github.com/hendrycks/robustness);
iWildCam models. We use the iWildCam models from the open source library (https://worksheets.codalab.org/worksheets/0x52cea64d1d3f4fa89de326b4e31aa50a) which offers models trained/fine-tuned on the iWildCam training set (Beery et al., 2020). The models utilized in our paper are listed below:
@@
Appendix F Additional Visualisations
F.1 Reliability Diagram
We show the reliability diagrams of various calibration methods and classification models for the in-distribution set in Fig. 7 and out-of-distribution (OOD) set in Fig. 8
Figure 7: Reliability diagram of various models on the ImageNet validation set (25 bins). Bars above the dashed line indicate underconfidence, while those below indicate overconfidence. Our method effectively mitigates both overconfident and underconfident predictions across different scenarios.
Figure 8: Reliability diagram of various models on the out-of-distribution (OOD) ObjectNet dataset (25 bins). Bars above the dashed line indicate underconfidence, while those below indicate overconfidence. ObjectNet is a highly challenging OOD test set, where models often exhibit severe overconfidence. Our method significantly mitigates this issue.
F.2 AUROC curves